what is the density of amonia at 27 degree celcius and 3 atm pressure

Grade:12th pass

1 Answers

nino chan
2 years ago
PV=nRT 
P=3atm
T=27c=300k
n(moles)=given mass/molecular mass
PV=m/M*RT
PM/RT=m/v (m/v=density)
3*molecular mass of ammonia/8.31*300=density
3*27/8.31*300=density
do the math kid ^